# publications.bib

### v3

@@ 4.3. Verification of additional assumptions
where we have used that localization commutes with direct sums since it is a left adjoint by definition. Now, using the compatibility of the Fargues-Scholze correspondence with central isogenies [FS, Theorem IX.6.1], either $\phi^{\prime}=g\circ\phi$ and $A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})}\in\mathrm{D}(G^{\prime}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}$ and, by the idempotence of the localization map, we have that $(A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}=A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})}$ or $(A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}$ is 00. The claim follows. ∎
@@
By the previous lemma applied to $G^{\mathrm{der}}=G^{\prime}\subset G$, we have a natural decomposition
@@
Taking limits over $K^{p}$, we obtain
$H^{i}(\mathrm{Sh}(\mathbf{G},X),\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}\simeq\bigoplus_{\begin{subarray}{c}\phi\\ \phi^{\prime}=g\circ\phi\end{subarray}}H^{i}(\mathrm{Sh}(\mathbf{G},X),\overline{\mathbb{F}}_{\ell})_{\phi}.$
Hence, we see that $H^{i}(\mathrm{Sh}(\mathbf{G},X),\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}$ vanishes for $i<d$, since all the $\phi$ appearing on the right-hand side are generic, and hence we can apply Theorem 1.16 and take limits to see that all the direct summands vanish. Applying $R\Gamma(K_{p},-)$ and Lemma 4.2 (3), we see that
$H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$
vanishes for $i<d$. Thus, we see that $H^{i}(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$ vanishes for $i<d$, and therefore the same is true for $H^{i}(\mathrm{Sh}(\mathbf{G}_{2},X_{2})_{K_{2,p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ from the discussion above. By the Hochschild-Serre spectral sequence, we see that for all sufficiently small $K_{2}^{p}$, $H^{i}_{c}(\mathrm{Sh}(\mathbf{G}_{2},X_{2})_{K_{2,p}K_{2}^{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ vanishes for $i<d$.
Now, consider a generic maximal ideal $\mathfrak{m}_{2}$ for the spherical Hecke algebra $H_{K_{2,p}}$ of $G_{2}$. This corresponds to a generic maximal ideal $\mathfrak{m}^{\prime}$ of $H^{\prime}_{K^{\prime}_{p}}$. It remains to observe that for any finitely-generated $H_{K_{2,p}}$-module $A$, if the localization $A_{\mathfrak{m}^{\prime}}=0$, then there is some element in $r\in H^{\prime}_{K^{\prime}_{p}}\backslash\mathfrak{m}^{\prime}$ such that $rA=0$. Thus we must have $A_{\mathfrak{m}_{2}}=0$ as well since $H^{\prime}_{K^{\prime}_{p}}\backslash\mathfrak{m}^{\prime}\subset H_{K_{2,p}}\backslash\mathfrak{m}_{2}$. ∎
As a corollary, we can strengthen previous results of Caraiani-Tamiozzo [CT, Theorem B], who previously showed torsion vanishing for Hilbert modular varieties under the additional assumption that $p$ was split in the totally real field $F$ (though we also note that they showed torsion vanishing under a hypothesis on $\mathfrak{m}$ which is weaker than the genericity considered here, see Remark 1.7).
@@
Conjecture 1.2 is true for Hilbert-Siegel Shimura varieties (attached to $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GL}_{2}$, $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$) and quaternionic Shimura varieties.
@@
Observe that for Hilbert-Siegel Shimura varieties (attached to $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GL}_{2}$, $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$), there is a cover by a PEL-type Shimura variety with local group $G$ of the form $G(\mathrm{SL}_{2})$ and $G(\mathrm{Sp}_{4})$ respectively. For the case of quaternionic Shimura varieties, we can relate their geometric connected components to unitary PEL-type Shimura varieties with local group $\mathrm{GU}_{2}$, as described in [TX16, Corollary 3.11]. Therefore, the result follows from Theorem 1.16. ∎

### v5

@@ 5. The Proof of the Main Theorems
where we have used that localization commutes with direct sums since it is a left adjoint by definition. Now, using the compatibility of the Fargues-Scholze correspondence with central isogenies [16, Theorem IX.6.1], either $\phi^{\prime}=g\circ\phi$ and $A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})}\in\mathrm{D}(G^{\prime}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}$ and, by the idempotence of the localization map, we have that $(A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}=A_{\phi^{\prime}}|_{G^{\prime}(\mathbb{Q}_{p})}$ or $(A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}$ is 0. The claim follows. ∎
This passage to derived subgroups allows us to pass to connected components of Shimura varieties. To see this, fix a connected component $X^{+}\subset X$. This also fixes a $X^{+}_{1}\subset X_{1}$, and an isomorphism $X^{+}\simeq X^{+}_{1}$, since $\mathbf{G},\mathbf{G}_{1}$ have isomorphic adjoint quotients. For any compact open subgroup $K\subset\mathbf{G}(\mathbb{A}_{f})$, we let $\mathrm{Sh}^{+}(\mathbf{G},X)_{K}$ be the geometrically connected component which is the image of $X^{+}\times 1$. Moreover, we will let
Note that since all the transition morphisms are finite étale and hence affine, $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}$ is also a qcqs scheme by [61, Lemma 01YX].
Now, observe that the action of $H^{\prime}_{K^{\prime}_{p}}$ on $\mathrm{Sh}(\mathbf{G},X)_{K_{p}}$ via the inclusion $H^{\prime}_{K^{\prime}_{p}}\subset H_{K_{p}}$ preserves the geometric connected component, as it acts trivially on the set of geometric connected components. To see this, recall from [14, §2.1.3] that the connected components are given by the profinite group
$\pi_{0}(\mathrm{Sh}_{K_{p}}(\mathbf{G},X))=\mathbf{G}(\mathbb{Q})^{-}_{+}\backslash\mathbf{G}(\mathbb{A}^{f})/K_{p},$
where as in the notation of [14] we denote $\mathbf{G}(\mathbb{R})_{+}$ the inverse image of $\mathbf{G}^{\mathrm{ad}}(\mathbb{R})$, $\mathbf{G}(\mathbb{Q})_{+}=\mathbf{G}(\mathbb{R})_{+}\cap\mathbf{G}(\mathbb{Q}$ and $\mathbf{G}(\mathbb{Q})^{-}_{+}$ is the closure of $\mathbf{G}(\mathbb{Q})_{+}$ in $\mathbf{G}(\mathbb{A}_{f})$. Since this identification is $\mathbf{G}(\mathbb{A}_{f})$-equivariant, and $\pi_{0}(\mathrm{Sh}_{K_{p}}(\mathbf{G},X))$ is abelian, we thus we see that the action of $\mathbf{G}(\mathbb{A}^{f})$ must factor through the maximal abelian quotient, so the action of $\mathbf{G}^{\mathrm{der}}(\mathbb{A}^{f})$ is trivial.
Now, using the notation of [36, §3.3] we see that there exist groups $\mathscr{A}(\mathbf{G})^{\circ},\mathscr{A}(\mathbf{G}_{1}),\mathscr{A}(\mathbf{G})$, whose definition we do not precisely recall here, such that from [36, Proposition 3.3.10], we know that we have an isomorphism
$\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}}\simeq(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}\times\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ}.$
We further note that since $G_{1}^{\mathrm{der}}=G^{\mathrm{der}}$, we can simplify this to
$\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}}\simeq\mathrm{Sh}(\mathbf{G},X)_{K_{p}}^{+}\times\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ},$
since from [36, §3.3.9] we see that the map $\mathscr{A}(\mathbf{G})^{\circ}\simeq\mathscr{A}(\mathbf{G}^{\mathrm{der}})^{\circ}\simeq\mathscr{A}(\mathbf{G}^{\mathrm{der}}_{1})^{\circ}\hookrightarrow\mathscr{A}(\mathbf{G}_{1}))$ is injective, using that $\mathbf{G},\mathbf{G}_{1}$ have isomorphic derived subgroups.
Thus, we have an isomorphism
$R\Gamma(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}\simeq R\Gamma(\mathrm{Sh}(\mathbf{G},X)^{+}_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}\otimes^{\mathbb{L}}_{\overline{\mathbb{F}}_{\ell}}C(\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ})$
where $C(\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ})$ denotes the space of all continuous $\overline{\mathbb{F}}_{\ell}$-valued functions on the profinite set $\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ}$.
Thus, observe that if $R\Gamma(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$ is concentrated in degrees $[d,2d]$, then so is $R\Gamma(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$, where we note that the above tensor product has no higher derived terms. Summarizing the above discussion, we thus have:
Proposition 5.4.
Suppose that we have two Shimura data $(\mathbf{G},X)$, $(\mathbf{G}_{1},X_{1})$ such that $(\mathbf{G}^{\mathrm{ad}},X^{\mathrm{ad}})\simeq(\mathbf{G}^{\mathrm{ad}}_{1},X^{\mathrm{ad}}_{1})$, $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$, and both $G,G_{1}$ are unramified. Then the action of $H^{\prime}_{K^{\prime}_{p}}\subset H_{K^{\mathrm{hs}}_{p}}$ stabilizes each connected component, and for any maximal ideal $\mathfrak{m}^{\prime}\subset H^{\prime}_{K^{\prime}_{p}}$ if $R\Gamma(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$ is concentrated in degrees $[d,2d]$, then so is $R\Gamma(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$.
Proposition 5.5.
Suppose that $(\mathbf{G},X)$ is a PEL-type Shimura datum such that Conjecture 1.2 holds for $(\mathbf{G},X)$, and we have another Shimura datum $(\mathbf{G},X)$ such that $(\mathbf{G}^{\mathrm{ad}},X^{\mathrm{ad}})\simeq(\mathbf{G}^{\mathrm{ad}}_{1},X^{\mathrm{ad}}_{1})$, $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$, and both $G,G_{1}$ are unramified. Then Conjecture 1.2 holds for $(\mathbf{G}_{1},X_{1})$.
Proof.
Since we assume that Conjecture 1.2 is true for $(\mathbf{G},X)$, we will first show that a maximal ideal $\mathfrak{m}$ of $H_{K_{p}}$ is generic if and only if the maximal ideal $\mathfrak{m}^{\prime}$ of $H^{\prime}_{K^{\prime}_{p}}$ is generic. To see this, we will reformulate this in terms of $L$-parameters. This is equivalent to showing that an $L$-parameter
$\phi=\phi_{\mathfrak{m}}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$
attached to $\mathfrak{m}$ is generic if and only if the composition $\phi^{\prime}$ with the map $g:\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})\rightarrow\phantom{}^{L}T^{\prime}(\overline{\mathbb{F}}_{\ell})$ induced by the inclusion of tori $T^{\prime}\hookrightarrow T$ is generic. (Here, $T^{\prime}=G^{\mathrm{der}}\cap T$). This follows from the observation that any coroot $\alpha$ factors through $G^{\mathrm{der}}$, and hence the composition $\alpha\circ\phi$ is equal to $\alpha\circ\phi^{\prime}$.
@@
By Lemma 5.3 applied to $G^{\mathrm{der}}=G^{\prime}\subset G$, we have a natural decomposition
@@
Hence, we see that $H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K^{p}},\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}$ vanishes for $i<d$, since all the $\phi$ appearing on the right-hand side are generic, and hence we can apply Theorem 1.8. Applying $R\Gamma(K_{p}^{\prime},-)$ and Lemma 4.3 (3), we see that
$H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K_{p}^{\prime}K^{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$
vanishes for $i<d$. Now, we can further apply the Hochschild-Serre spectral sequence to the $K_{p}/K_{p}^{\prime}$-torsor $\mathrm{Sh}(\mathbf{G},X)_{K_{p}^{\prime}K^{p}}\rightarrow\mathrm{Sh}(\mathbf{G},X)_{K_{p}K^{p}}$ to get that $H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K_{p}K^{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ vanishes for $i<d$.
Now, we can consider the connected component $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}K^{p}}$, and observe that $\mathrm{Sh}(\mathbf{G},X)_{K_{p}K^{p}}$ is a finite disjoint union of isomorphic copies of $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}K^{p}}$. Moreover, exactly as explained above the action of $H^{\prime}_{K_{p}^{\prime}}$ stabilizes each connected component, so we see that $H^{i}(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}K^{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ vanishes for $i<d$ as well.
Finally, taking colimits over $K^{p}$, we see that $H^{i}(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ vanishes for $i<d$ as well. Applying Proposition 5.4, $H^{i}(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$ vanishes for $i<d$.
Now, we let $\mathbf{Z}\subset\mathbf{G}$ be the center, and we denote by $\mathbf{Z}(\mathbb{Q})^{-}$ the closure of $\mathbf{Z}(\mathbb{Q})$ in $\mathbf{Z}(\mathbb{A}_{f})$. Denote by $\overline{K_{1}^{p}}$ the quotient
$\overline{K_{1}^{p}}=(K_{1p}K^{p}_{1}/K_{1p}K^{p}_{1}\cap\mathbf{Z}(\mathbb{Q})^{-})/(K^{p}_{1}/K^{p}_{1}\cap\mathbf{Z}(\mathbb{Q})^{-}).$
By the Hochschild-Serre spectral sequence (applied to the quotients $\overline{K_{1}^{p}}$), we see that for all sufficiently small $K_{1}^{p}$, $H^{i}(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}K_{1}^{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ vanishes for $i<d$.
Now, consider a generic maximal ideal $\mathfrak{m}_{1}$ for the spherical Hecke algebra $H_{K_{1p}}$ of $G_{1}$. This corresponds to a generic maximal ideal $\mathfrak{m}^{\prime}$ of $H^{\prime}_{K^{\prime}_{p}}$. It remains to observe that for any finitely-generated $H_{K_{1p}}$-module $A$, if the localization $A_{\mathfrak{m}^{\prime}}=0$, then there is some element in $r\in H^{\prime}_{K^{\prime}_{p}}\backslash\mathfrak{m}^{\prime}$ such that $rA=0$. Thus we must have $A_{\mathfrak{m}_{1}}=0$ as well since $H^{\prime}_{K^{\prime}_{p}}\backslash\mathfrak{m}^{\prime}\subset H_{K_{1p}}\backslash\mathfrak{m}_{1}$. Taking $A=H^{i}(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}K_{1}^{p}},\overline{\mathbb{F}}_{\ell})$, we can conclude. ∎
We obtain the following by combining this Proposition with Theorem 1.8.
@@
Suppose $(\mathbf{G}_{1},X_{1})$ is an abelian-type Shimura datum which has an associated PEL-type datum $(\mathbf{G},X)$ of type AC satisfying Assumption 1.13 and such that $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$, $G_{1}:=\mathbf{G}_{1,\mathbb{Q}_{p}}$ is unramified and $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. We assume that $\mathfrak{m}_{1}\subset H_{K_{1p}^{\mathrm{hs}}}$ is a generic maximal ideal. Then, for a level $K_{1}=K_{1}^{p}K_{1p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, the cohomology of $R\Gamma(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1},\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ (resp. $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1},\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$) is concentrated in degrees $d\leq i\leq 2d$ (resp. $0\leq i\leq d$).
@@
This follows from Theorem 1.8, once we show that the generic condition on $\phi^{T}_{\mathfrak{m}}$ is also satisfied by $\phi^{T_{1}}_{\mathfrak{m}_{1}}$. To see this, we can in fact see that this condition is satisfied if and only if it is satisfied by $G^{\mathrm{der}}$, where we look at the composition
$\phi^{T^{\mathrm{der}}}_{\mathfrak{m^{\prime}}}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})\rightarrow\phantom{}^{L}T^{\mathrm{der}}(\overline{\mathbb{F}}_{\ell}).$
In fact, observe that we can descend any coroot $\alpha$ to a coroot $\alpha^{\mathrm{der}}:\phantom{}^{L}T^{\mathrm{der}}(\Lambda)\rightarrow\overline{\mathbb{F}}_{\ell}^{\times}$, since $G,G^{\mathrm{der}}$ have the same (co-)root systems, and we have $\alpha\circ\phi_{\mathfrak{m}}^{T}=\alpha^{\mathrm{der}}\circ\phi^{T^{\mathrm{der}}}_{\mathfrak{m^{\prime}}}$. ∎
In particular, we can strengthen previous results of Caraiani-Tamiozzo [9, Theorem B], who previously showed torsion vanishing for Hilbert modular varieties under the additional assumption that $p$ was split in the totally real field $F$ (though we also note that they showed torsion vanishing under a hypothesis on $\mathfrak{m}$ which is weaker than the genericity considered here, see Remark 1.7).
Corollary 5.7.
Conjecture 1.2 is true for Hilbert modular varieties, quaternionic Shimura varieties assuming $(\ell,2[L:\mathbb{Q}_{p}])=1$, and Hilbert-Siegel Shimura varieties (attached to $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$) assuming that $\ell\nmid[L:\mathbb{Q}_{p}]$, where for both cases $L$ is the completion of $F$ at some unramified prime above $p$.
Proof.
Observe that for Hilbert-Siegel Shimura varieties (attached to $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GL}_{2}$, $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$), there is a cover by a PEL-type Shimura variety with local group $G$ of the form $G(\mathrm{SL}_{2})$ and $G(\mathrm{Sp}_{4})$ respectively. For the case of quaternionic Shimura varieties, we can relate their geometric connected components to unitary PEL-type Shimura varieties with local group $G=\mathrm{G}(U_{2}\times\dots\times U_{2})$, as described in [64, Corollary 3.11]. This group does not appear in Table 1, but we show below that it satisfies the additional assumptions to apply Theorem 4.23. Therefore, the result follows from Corollary 5.6. ∎
Lemma 5.8.
The group $G=\mathrm{G}(U_{2}\times\dots\times U_{2})$ satisfies Assumption 4.5, and genericity implies regularity and strong $\mu$-regularity for any toral parameter.
Proof.
Firstly observe that $G$ is a subgroup of $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GU}_{2}$. Thus, we can apply Proposition 4.9 because we know the result for $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GU}_{2}$. The additional assumptions required to invoke Theorem 4.23 can be verified after passing to an extension where the group splits, from which we see that $\mathrm{G}(U_{2}\times\dots\times U_{2})$ is isomorphic to $G^{\prime}=(\mathbb{G}_{m}\times\mathbb{G}_{m}\times\mathrm{GL}_{2}\times\dots\times\mathrm{GL}_{2})/\mathbb{G}_{m}$, where the $\mathbb{G}_{m}$ in the quotient is the diagonal copy of $\mathbb{G}_{m}$. Moreover, we observe that the maximal torus $T^{\prime}$ of $G^{\prime}$ can be further identified with $\mathbb{G}_{m}\times T\times\dots\times T$ by setting the element in the first copy of $\mathbb{G}_{m}$ to 1, and where each $T$ is the diagonal torus on $\mathrm{GL}_{2}$. In particular, we can reduce to the case of a product of $\mathrm{GL}_{2}$’s, from which the additional conditions follow as in the proof of Lemma 4.25. More precisely, we write an element of $T$ as $(t,t_{11},t_{12},\dots,t_{d1},t_{d2})$, and the character $\chi$ as $\nu(t)\chi_{11}(t_{11})\dots\chi_{d2}(t_{d2})$. We get regularity by observing that if any element of $W_{G}$ is non-trivial on some $T$ factor (say the $i$th one), then by setting $t=1$, $t_{j1}=t_{j2}=1$, we get the relation $\chi_{i1}\chi^{-1}_{i2}\simeq\mathbf{1}$, a contradiction. We get $\mu$-regularity for any dominant cocharacter $\mu$, since we see that the strong $\mu$-regularity condition again holds for any minuscule $\mu$, and any dominant cocharacter of $G$ can be written as a sum of such minuscule $\mu$. ∎
Remark 5.9.
As remarked by a referee, there are Shimura data other than the ones listed in Corollary 5.7 to which we can apply Proposition 5.5. For instance, one can consider the Shimura data where the group $\mathbf{G}_{1}$ is an inner form of $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$ for $F$ totally real, and which is split at all real places, and $X_{1}$ is the product of $F:\mathbb{Q}]$ copies of the Siegel upper half space $\mathcal{H}_{2}$. In this case, one can show that there exists a PEL-type datum $(\mathbf{G},X)$ satisfying $(\mathbf{G}^{\mathrm{ad}},X^{\mathrm{ad}})\simeq(\mathbf{G}^{\mathrm{ad}}_{1},X^{\mathrm{ad}})$, and $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$.